40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Sep 19 01:49:52 2020, Koji, Update, General, M4.5 EQ in LA 
    Reply  Sat Sep 19 11:27:04 2020, rana, Update, General, M4.5 EQ in LA Sep18-EQ.pdf
       Reply  Sat Sep 19 18:08:34 2020, rana, Update, General, M4.5 EQ in LA 
    Reply  Sat Sep 19 19:14:59 2020, Koji, Update, General, ITMX released / ETMY UR magnet knocked off IMG_6218.jpeg
       Reply  Sun Sep 20 11:41:54 2020, gautam, Update, General, IMC re-locked 
       Reply  Mon Sep 21 12:40:58 2020, gautam, Update, General, Workable IFO recovery IFOrecovery.pngIMG_5345.JPG
Message ID: 15585     Entry time: Sat Sep 19 19:14:59 2020     In reply to: 15580     Reply to this: 15588   15590
Author: Koji 
Type: Update 
Category: General 
Subject: ITMX released / ETMY UR magnet knocked off 

There were two SUSs which didn't look normal.

- ITMX was easily released by the bias slider -> Shake the pitch slider and while all the OSEM values are moving, turn on the damping control (with x10 large watchdog threshold)

- ETMY has UR OSEM 0V output. This means that there is no light. And this didn't change at all with the slider move.
- Went to the Y table and tried to look at the coils. It seems that the UR magnet is detached from the optic and stuck in the OSEM.

We need a vent to fix the suspension, but until then what we can do is to redistribute the POS/PIT/YAW actuations to the three coils.

Attachment 1: IMG_6218.jpeg  453 kB  Uploaded Sat Sep 19 20:47:32 2020  | Hide | Hide all
IMG_6218.jpeg
ELOG V3.1.3-